# State-observation augmented diffusion model for nonlinear assimilation with unknown dynamics

### v2

@@ 3.3.1 Augmented dynamical system
$\bm{y}_{k}=\mathcal{H}_{k}(\bm{x}_{k})+\bm{\varepsilon}_{k}=\bm{S}_{k}\mathcal{H}(\bm{x}_{k})+\bm{\varepsilon}_{k}=\bm{T}_{k}\bm{z}_{k}+\bm{\varepsilon}_{k},\quad\bm{\varepsilon}_{k}\sim\mathscr{D}^{Y}.$ (26)
@@
$\begin{cases}\bm{z}_{k}=\mathcal{M}_{k}^{Z}\left(\bm{z}_{k-1},\bm{\eta}_{k}\right),&\bm{\eta}_{k}\sim\mathscr{D}^{X},\\ \bm{y}_{k}=\bm{T}_{k}\bm{z}_{k}+\bm{\varepsilon}_{k},&\bm{\varepsilon}_{k}\sim\mathscr{D}^{Y}.\end{cases}$ (27)
@@
$\bm{y}_{S:T}=\bm{T}_{S:T}\bm{z}_{S:T}+\bm{\varepsilon}_{S:T},\quad\bm{\varepsilon}_{S:T}\sim\mathscr{D}_{S:T}^{Y}$ (28)
@@ 3.3.3 Forward-diffusion corrector
In our augmented model (27), the observation $\bm{y}$ is a linear transformation of $\bm{z}$, or more specifically, a subsampled version of $\bm{z}$ with observational noise as described in eq. 26. To stabilize the reverse-time diffusion process, we propose directly replacing part of the denoised results with a perturbed version of $\bm{y}$. More precisely, let $\bm{z}_{t}$ be the diffusion state at time $t$ with an initial state $\bm{z}_{0}=\bm{z}$, then
@@
whenever $\sigma_{t}^{2}\bm{I}-\mu_{t}^{2}\bm{\Sigma}^{Y}$ is positive semidefinite.
@@ 3.3.4 Implementation
21:function ForwardDiffusionCorrector($\bm{z}_{t}$, $t$)
@@ 4.2 Network structures and training procedures
To ensure a fair comparison with the SDA baseline, we adopt the same network structure for the denoising network $\bm{\varepsilon}_{\theta}(\bm{x}_{t},t)$, which utilizes a U-Net structure mixed with temporal convolutions and embeddings for diffusion time steps $t$. We have included the detailed architecture in appendix B for completeness. The only difference between our network and the SDA baseline lies in the input channels. SDA learns a generative model for the state $\bm{x}_{S:T}$, while our SOAD learns for the augmented state $\bm{z}_{S:T}$. Since each augmented state includes a clean observation $\mathcal{H}(\bm{x})_{S:T}$, training SOAD requires additional calls of the observation operator $\mathcal{H}$.
@@
$L(\bm{\theta})=\frac{1}{2J}\sum_{j=1}^{J}\left\|\bm{\varepsilon}_{\bm{\theta}}\left(\mu_{t_{j}}\bm{z}_{S_{j}:S_{j}+8}^{(j)}+\sigma_{t_{j}}\bm{\varepsilon}_{j},t_{j}\right)-\bm{\varepsilon}_{j}\right\|^{2}$ (49)
@@ 4.3 Observations
$\bm{y}_{k}=\bm{S}_{k}\circ\mathcal{H}(\bm{x}_{k})+\bm{\varepsilon}_{k},\quad\bm{\varepsilon}_{k}\sim\mathcal{N}\left(0,0.1^{2}\bm{I}\right),$ (51)
@@
Readers may refer to fig. 4 for detailed illustrations. Throughout our experiments, we evaluate random ratios $p\in\{1.0,0.25,0.625,0.01\}$ and strides $s\in\{1,2,4,10\}$ for partial observations under varying time intervals $N\in\{1,2,4,8\}$.
@@ 4.5 Assimilation experiments
We consider the following three assimilation settings.
@@ 4.5.1 Assimilation with background prior
$\bm{y}_{0}=\bm{x}_{0}+\bm{\varepsilon}_{0},\quad\bm{\varepsilon}_{0}\sim\mathcal{N}\left(0,(\sigma^{\mathrm{b}})^{2}\bm{I}\right),$ (56)
@@
Our experiments start with the element-wise observations $\mathcal{H}^{\mathrm{e}}$ and $\mathcal{H}^{\mathrm{h}}$. The averaged RMSEs over all the 9 time steps for the assimilated states are summarized in fig. 5.
@@
(b) SDA (baseline) for $\mathcal{H}^{\mathrm{e}}$
(c) SOAD (ours) for $\mathcal{H}^{\mathrm{h}}$
(d) SDA (baseline) for $\mathcal{H}^{\mathrm{h}}$
@@
Our SOAD model performs slightly worse as the time interval $N$ increases and the observational ratio $p$ or $(1/s)^{2}$ decreases, which aligns with our intuition that the assimilation task becomes more difficult with fewer observations. In the most challenging case, where the observations are available only every 8 time steps with a ratio of $1\%$, the RMSEs are around 0.2 and 0.3 for the easy and hard observations, respectively. This indicates that our method is still able to extract the physical features from rare observations to some extent. In contrast, the SDA baseline shows a significant performance drop even when the observations are available everywhere for each time step. The inferior performance is likely due to the linearity assumption for the observation operator in the SDA formulations, which is unsuitable for highly nonlinear observations.
Additionally, the step-wise RMSEs for our SOAD model over the assimilation window are exhibited in fig. 6. In both observation cases, The increments of RMSEs at observational ratios of 100% and 25% are mild, and no significant changes in RMSEs are observed when the easier $\mathcal{H}^{\mathrm{e}}$ is replaced with the harder $\mathcal{H}^{\mathrm{h}}$. Moreover, relatively lower assimilation errors are obtained whenever observational data are available, supporting the idea that more observations enhance the assimilation process. By comparing results with the same observational ratios ($p=(1/s)^{2}$), we may conclude that regular (uniform) observations are more beneficial for the assimilation process than irregular (random) ones.
@@
(b) the hard observation $\mathcal{H}^{\mathrm{h}}$
Figure 6: Step-wise RMSEs of the assimilated states via our SOAD approach.
To explore the potential of our SOAD model in real applications, we also test it with the vorticity-to-velocity mapping $\mathcal{H}^{\mathsf{v2v}}$, and the results are shown in fig. 7. We do not include the performance of the SDA baseline since it diverges in all cases. The averaged RMSEs in fig. 7(a) are lower than those in figs. 5(a) and 5(c), which may be attributed to the linearity of the observational operator $\mathcal{H}^{\mathsf{v2v}}$. The results indicate that our SOAD approach is capable of handling the vorticity-to-velocity mapping, a more practical observational model in real applications. Although SOAD shows faster increasing RMSEs with higher uncertainty for $\mathcal{H}^{\mathsf{v2v}}$ (fig. 7(b)) compared with $\mathcal{H}^{\mathrm{e}}$ and $\mathcal{H}^{\mathrm{h}}$ (fig. 6), the subsequent experiments may offer some evidences that the assimilation process is unlikely to diverge.
(a) averaged RMSEs for SOAD with $\mathcal{H}^{\mathsf{v2v}}$
(b) step-wise RMSEs for SOAD with $\mathcal{H}^{\mathsf{v2v}}$
Figure 7: Performances of our SOAD for assimilation with prior when the observational model is $\mathcal{H}^{\mathsf{v2v}}$.
In practice, we cannot always expect the availability of the exact probability distribution for background prior. Background errors may result from many aspects, such as unresolved sub-scale features, temporal interpolations and physical parameterizations. Therefore, we investigate whether the discrepancy between our background prior model (56) and the real distribution affects the assimilation performance. We conduct experiments with the same settings as before but feed a background prior with unknown noise
into the model instead of $\bm{y}_{0}$. Here, $\mathcal{M}^{\mathrm{QG}}$ is the forward propagation model of the QG equations, and $\bm{x}_{-1}$ stands for the snapshot at the previous time step. Comparisons between the step-wise RMSEs are displayed in fig. 8, where we fix $N=1$ and $p=0.25$. Even with incorrect background prior assumptions, our SOAD model has the ability of self-correction with a few time steps of observations. Another interesting observation is that the RMSEs for the hard observation $\mathcal{H}^{\mathrm{h}}$ are lower than those for $\mathcal{H}^{\mathrm{e}}$ and $\mathcal{H}^{\mathsf{v2v}}$. One possible explanation is that when the observational data is sufficient for the model to capture the dynamics, observations from a harder operator $\mathcal{H}^{\mathrm{h}}$ may provide more detailed information since $\mathcal{H}^{\mathrm{h}}$ is much more sensitive to the input.
@@ 4.5.2 Assimilation without background prior
All assimilation results are recorded as heatmaps of RMSEs in fig. 9. In general, our SOAD model performs well for both the $\mathcal{H}^{\mathrm{e}}$ and the $\mathcal{H}^{\mathsf{v2v}}$ observations, with slightly higher accuracy for the latter. The observation $\mathcal{H}^{\mathrm{h}}$ appears to be the most challenging across all the testing observational operators.
To visualize the assimilated states, we have plotted the assimilated vorticity with random observational ratios $p\in\{0.25,0.0625,0.01\}$ in figs. 10, 11 and 12 for $\mathcal{H}^{\mathrm{e}}$, $\mathcal{H}^{\mathrm{h}}$ and $\mathcal{H}^{\mathsf{v2v}}$, respectively. For clarity, we only include the vorticity and the associated observations for the first layer, as they contain more small-scale details. For the $\mathcal{H}^{\mathrm{e}}$ observation, our SOAD is capable of recovering most of the physical features until $p$ reaches $1\%$, in which case the assimilated states still share much similarity with the ground truth. With the harder $\mathcal{H}^{\mathrm{h}}$, the SOAD model fails to reconstruct the real system states when $p\leq 0.0625=(1/4)^{2}$, consistent with the previous results shown in fig. 9(b). Surprisingly, the assimilated states with rare observations do not collapse and still follow certain dynamics, suggesting that the lack of observational data, rather than incapacity of learning the dynamics, leads to the failure. Finally, when the observation is the vorticity-to-velocity mapping $\mathcal{H}^{\mathsf{v2v}}$, the assimilated states are closer to the ground truth even when observations are available at only $1\%$ of the grids. We believe that the success may result from the linearity of $\mathcal{H}^{\mathsf{v2v}}$. The results indicate that our SOAD approach is stable when handling highly nonlinear observations in the long term performs well on non-element-wise observations such as the vorticity-to-velocity mapping $\mathcal{H}^{\mathsf{v2v}}$, which is a promising sign for real applications.
@@ 4.5.3 Assimilation with multiple observations
(a) averaged RMSEs for collections containing $\mathcal{H}^{e}$
(b) averaged RMSEs for collections containing $\mathcal{H}^{h}$
@@
Figure 13: RMSEs for multi-modal assimilation with our SOAD model. The error uncertainty introduced by different random seeds are marked on the top of each bar. Duplicate results may appear in multiple subplots for illustration purposes.

### v3

@@ 3.3.1 Augmented dynamical system
$\bm{y}_{k}=\mathcal{H}_{k}(\bm{x}_{k})+{\bm{\epsilon}_{k}}=\bm{S}_{k}\mathcal{H}(\bm{x}_{k})+{\bm{\epsilon}_{k}}=\bm{T}_{k}\bm{z}_{k}+{\bm{\epsilon}_{k}},\quad{\bm{\epsilon}_{k}}\sim\mathscr{D}^{Y}.$ (26)
@@
$\begin{cases}\bm{z}_{k}=\mathcal{M}_{k}^{Z}\left(\bm{z}_{k-1},\bm{\eta}_{k}\right),&\bm{\eta}_{k}\sim\mathscr{D}^{X},\\ \bm{y}_{k}=\bm{T}_{k}\bm{z}_{k}+{\bm{\epsilon}_{k}},&{\bm{\epsilon}_{k}}\sim\mathscr{D}^{Y}.\end{cases}$ (27)
@@
$\bm{y}_{S:T}=\bm{T}_{S:T}\bm{z}_{S:T}+{\bm{\epsilon}_{S:T}},\quad{\bm{\epsilon}_{S:T}}\sim\mathscr{D}_{S:T}^{Y}$ (28)
@@ 3.3.3 Forward-diffusion corrector
In our augmented model (27), the observation $\bm{y}$ is a linear transformation of $\bm{z}$, or more specifically, a subsampled version of $\bm{z}$ with observational noise as described in eq. 26. To stabilize the reverse-time diffusion process, we propose directly replacing part of the denoised results with a perturbed version of $\bm{y}$. The correction is based on an auxiliary forward diffusion of the observational data, which is then used to guide or adjust the denoising states where observations are available. Consequently, we refer to such a corrector as the “forward-diffusion corrector”.
More precisely, let $\bm{z}_{t}$ be the diffusion state at time $t$ with an initial state $\bm{z}_{0}=\bm{z}$, then
@@
whenever the matrix $\left(\sigma_{t}^{2}\bm{I}-\mu_{t}^{2}\bm{\Sigma}^{Y}\right)$ is positive semidefinite.
@@ 3.3.4 Implementation
21:function FDC($\bm{z}_{t}$, $t$$\triangleright$ the forward-diffusion corrector
@@ 4.2 Network structures and training procedures
To ensure a fair comparison with the SDA baseline, we adopt the same network structure for the denoising network $\bm{\varepsilon}_{\theta}(\bm{x}_{t},t)$, which utilizes a U-Net structure mixed with temporal convolutions and embeddings for diffusion time steps $t$. We have included the detailed architecture in appendix B for completeness. The only difference between our network and the SDA baseline lies in the input channels. SDA learns a generative model for the state $\bm{x}_{S:T}$, while our SOAD learns for the augmented state $\bm{z}_{S:T}$, which means that the corresponding observations $\bm{y}_{S:T}$ is fed into the network for training as well. Since each augmented state includes a clean observation $\mathcal{H}(\bm{x})_{S:T}$, training SOAD requires additional calls of the observation operator $\mathcal{H}$.
@@
$\min_{\bm{\theta}}L(\bm{\theta})=\min_{\bm{\theta}}\frac{1}{2J}\sum_{j=1}^{J}\left\|\bm{\varepsilon}_{\bm{\theta}}\left(\mu_{t_{j}}\bm{z}_{S_{j}:S_{j}+8}^{(j)}+\sigma_{t_{j}}\bm{\varepsilon}_{j},t_{j}\right)-\bm{\varepsilon}_{j}\right\|^{2}$ (49)
@@ 4.3 Observations
$\bm{y}_{k}=\bm{S}_{k}\circ\mathcal{H}(\bm{x}_{k})+{\bm{\epsilon}_{k}},\quad{\bm{\epsilon}_{k}}\sim\mathcal{N}\left(0,0.1^{2}\bm{I}\right),$ (51)
@@
Readers may refer to fig. 4 for detailed illustrations, where we only visualize the first-layer vorticity ($q_{1}$) for simplicity. Throughout our experiments, we evaluate random ratios $p\in\{1.0,0.25,0.625,0.01\}$ and strides $s\in\{1,2,4,10\}$ for partial observations under varying time intervals $N\in\{1,2,4,8\}$.
@@ 4.5 Assimilation experiments
We consider the following four assimilation settings.
@@
  4. (IV)
    Assimilation with non-Gaussian distributions as an extension of setting (I) and (II), where we set the background prior and the observational noise as non-Gaussian to study the robustness of our model.
@@ 4.5.1 Assimilation with background prior
$\bm{y}_{0}=\bm{x}_{0}+{\bm{\epsilon}_{0}},\quad{\bm{\epsilon}_{0}}\sim\mathcal{N}\left(0,(\sigma^{\mathrm{b}})^{2}\bm{I}\right),$ (56)
@@
Our experiments start with the element-wise observations $\mathcal{H}^{\mathrm{e}}$ and $\mathcal{H}^{\mathrm{h}}$. Besides, to explore the potential of our SOAD model in real applications, we also test it with the vorticity-to-velocity mapping $\mathcal{H}^{\mathsf{v2v}}$. The averaged RMSEs over all the 9 time steps for the assimilated states are summarized in fig. 5.
@@
(b) SOAD (no f.d.c.) for $\mathcal{H}^{\mathrm{e}}$
(c) SDA (baseline) for $\mathcal{H}^{\mathrm{e}}$
(d) SOAD (ours) for $\mathcal{H}^{\mathrm{h}}$
(e) SOAD (no f.d.c.) for $\mathcal{H}^{\mathrm{h}}$
(f) SDA (baseline) for $\mathcal{H}^{\mathrm{h}}$
(g) SOAD (ours) for $\mathcal{H}^{\mathsf{v2v}}$
(h) SOAD (no f.d.c.) for $\mathcal{H}^{\mathsf{v2v}}$
(i) SDA (baseline) for $\mathcal{H}^{\mathsf{v2v}}$
@@
Our SOAD model performs slightly worse as the time interval $N$ increases and the observational ratio $p$ or $(1/s)^{2}$ decreases, which aligns with our intuition that the assimilation task becomes more difficult with fewer observations. In the most challenging case, where the observations are available only every 8 time steps with a ratio of $1\%$, the RMSEs are around 0.2 and 0.3 for the easy and hard observations, respectively. This indicates that our method is still able to extract the physical features from rare observations to some extent. Another interesting observation is that the RMSEs for the hard observation $\mathcal{H}^{\mathrm{h}}$ are lower than those for $\mathcal{H}^{\mathrm{e}}$ and $\mathcal{H}^{\mathsf{v2v}}$. One possible explanation is that when the observational data is sufficient for the model to capture the dynamics, observations from a harder operator $\mathcal{H}^{\mathrm{h}}$ may provide more detailed information since $\mathcal{H}^{\mathrm{h}}$ is much more sensitive to the input. In contrast, the SDA baseline (figs. 5(c), 5(f) and 5(i)) shows a significant performance drop even when the observations are available everywhere for each time step. The inferior performance is likely due to the linearity assumption for the observation operator in the SDA formulations, which is unsuitable for highly nonlinear observations.
Meanwhile, fig. 5 includes an ablation study of our forward-diffusion corrector, abbreviated as “f.d.c.” in the captions, as well. To study the effectiveness of our forward-diffusion corrector, we have kept all the experimental settings the same as those for the SOAD and the SDA methods, but removed all the updates by the forward-diffusion corrector (line 12 and 16 in algorithm 1). By comparing figs. 5(a), 5(d) and 5(g) with figs. 5(b), 5(e) and 5(h) respectively we can conclude that the forward-diffusion corrector has played an important role in improving the assimilation performance. For the observations $\mathcal{H}^{\mathrm{e}}$ and $\mathcal{H}^{\mathsf{v2v}}$, our SOAD model without the forward-diffusion corrector has shown a similar pattern as the normal SOAD, which is to say, the denser and the more frequent observations we have for assimilation, the higher accuracy we can obtain. For the much harder observation $\mathcal{H}^{\mathrm{h}}$, the forward-diffusion corrector has prevented divergence and thus exhibited its great potential to stablize the generative process.
Additionally, the step-wise RMSEs for our SOAD model over the assimilation window are exhibited in fig. 6. Note that we start from a background prior as an initial estimate, which introduces much more physical information than the subsequent observational data. Although observations are available for some of the subsequent time steps, the physical information encoded in the observations may not be sufficiently informative to fully correct the trajectory as it evolves, particularly in the early stages of assimilation. This can result in increasing deviation from the ground truth at later time steps, especially if the dynamical system exhibits sensitivity to initial conditions or chaotic behavior. In both observation cases, The increments of RMSEs at observational ratios of 100% and 25% are mild, and no significant changes in RMSEs are observed when the easier $\mathcal{H}^{\mathrm{e}}$ is replaced with the harder $\mathcal{H}^{\mathrm{h}}$. Moreover, relatively lower assimilation errors are obtained whenever observational data are available, supporting the idea that more observations enhance the assimilation process. By comparing results with the same observational ratios ($p=(1/s)^{2}$), we may conclude that regular (uniform) observations are more beneficial for the assimilation process than irregular (random) ones. In addition, although the step-wise RMSEs increase along with the time steps, and the uncertainty of $\mathcal{H}^{\mathsf{v2v}}$ (figs. 6(e) and 6(f)) is slightly higher compared with $\mathcal{H}^{\mathrm{e}}$ (figs. 8(a) and 8(b)) and $\mathcal{H}^{\mathrm{h}}$ (figs. 8(c) and 8(d)), the subsequent experiments may offer some evidences that the assimilation process is unlikely to diverge.
@@
(b) the easy observation $\mathcal{H}^{\mathrm{e}}$
(c) the hard observation $\mathcal{H}^{\mathrm{h}}$
(d) the hard observation $\mathcal{H}^{\mathrm{h}}$
(e) the vorticity-to-velocity $\mathcal{H}^{\mathsf{v2v}}$
(f) the vorticity-to-velocity $\mathcal{H}^{\mathsf{v2v}}$
Figure 6: Step-wise RMSEs of SOAD for assimilation with prior.
@@ 4.5.2 Assimilation without background prior
All assimilation results are recorded as heatmaps of RMSEs in fig. 7. In general, our SOAD model performs well for both the $\mathcal{H}^{\mathrm{e}}$ and $\mathcal{H}^{\mathsf{v2v}}$ observational operators, with slightly higher accuracy for the latter. The observation $\mathcal{H}^{\mathrm{h}}$ appears to be the most challenging across all the testing observational operators. More specifically, as also discussed in section 4.5.1, when observations are densely available, a condition that is uncommon in practice, the observation $\mathcal{H}^{\mathrm{h}}$ appears to be more informative and helpful than $\mathcal{H}^{\mathrm{e}}$ and $\mathcal{H}^{\mathsf{v2v}}$. However, under more typical conditions with partial observations, our SOAD model achieves the best performance with $\mathcal{H}^{\mathsf{v2v}}$, followed $\mathcal{H}^{\mathrm{e}}$, This may be attribute to the linearity of $\mathcal{H}^{\mathsf{v2v}}$, despite its non-local nature and greater computational complexity.
As before, we present the step-wise RMSEs of our SOAD model in fig. 8. Compared to the results in fig. 6, the performance degrades when assimilating without a background prior–an expected outcome given the absence of an initial estimate at the first time step. Notably, the RMSE remains relatively stable across time, with only slight increases observed at the beginning and end of the assimilation window. This pattern is consistent with the behavior of smoothing methods, which have limited observational support near the temporal boundaries. Furthermore, the individual RMSE profiles across different experimental settings demonstrate that the SOAD model maintains consistent accuracy over time, providing strong evidence of its long-term stability and resistance to divergence.
To visualize the assimilated states, we have plotted the assimilated vorticity with random observational ratios $p\in\{0.25,0.0625,0.01\}$ in figs. 9, 10 and 11 for $\mathcal{H}^{\mathrm{e}}$, $\mathcal{H}^{\mathrm{h}}$ and $\mathcal{H}^{\mathsf{v2v}}$, respectively. For clarity, we only include the vorticity and the associated observations for the first layer, as they contain more small-scale details. For the $\mathcal{H}^{\mathrm{e}}$ observation, our SOAD is capable of recovering most of the physical features until $p$ reaches $1\%$, in which case the assimilated states still share much similarity with the ground truth. With the harder $\mathcal{H}^{\mathrm{h}}$, the SOAD model fails to reconstruct the real system states when $p\leq 0.0625=(1/4)^{2}$, consistent with the previous results shown in fig. 7(b). Surprisingly, the assimilated states with rare observations do not collapse and still follow certain dynamics, suggesting that the lack of observational data, rather than incapacity of learning the dynamics, leads to the failure. Finally, when the observation is the vorticity-to-velocity mapping $\mathcal{H}^{\mathsf{v2v}}$, the assimilated states are closer to the ground truth even when observations are available at only $1\%$ of the grids. We believe that the success may result from the linearity of $\mathcal{H}^{\mathsf{v2v}}$. The results indicate that our SOAD approach is stable when handling highly nonlinear observations in the long term performs well on non-element-wise observations such as the vorticity-to-velocity mapping $\mathcal{H}^{\mathsf{v2v}}$, which is a promising sign for real applications.
@@
(a) the easy observation $\mathcal{H}^{\mathrm{e}}$
(b) the easy observation $\mathcal{H}^{\mathrm{e}}$
(c) the hard observation $\mathcal{H}^{\mathrm{h}}$
(d) the hard observation $\mathcal{H}^{\mathrm{h}}$
(e) the vorticity-to-velocity $\mathcal{H}^{\mathsf{v2v}}$
(f) the vorticity-to-velocity $\mathcal{H}^{\mathsf{v2v}}$
Figure 8: Step-wise RMSEs of SOAD for assimilation without prior.
@@ 4.5.3 Assimilation with multiple observations
(a) averaged RMSEs for collections containing $\mathcal{H}^{\mathrm{e}}$
(b) averaged RMSEs for collections containing $\mathcal{H}^{\mathrm{h}}$
@@
Figure 12: RMSEs for multi-modal assimilation with our SOAD model. The error uncertainty introduced by different random seeds is marked on the top of each bar. Duplicate results may appear in multiple subplots for illustration purposes.
@@
4.5.4 Model robustness beyond Gaussian assumptions
All the previous experiments have employed Gaussian assumptions for both the background estimate (when available) and the observational noise. However, this idealization rarely holds in realistic applications. Commonly, during online deployment, the background uncertainty is either unknown or poorly characterized, and the observational noise may exhibit heavy tails or other non-Gaussian features due to sensor limitations and dynamic environmental factors. To assess the robustness of our proposed model under these more challenging but realistic conditions, we conduct additional experiments in which the Gaussian assumptions are deliberately violated. These experiments serve to evaluate the model’s adaptability and reliability in the presence of distributional uncertainty.
In practice, we cannot always expect the availability of the exact probability distribution for background prior. Background errors may result from many aspects, such as unresolved sub-scale features, temporal interpolations and physical parameterizations. To explore the impact of potential mismatch between our assumed background prior model (56) and the true underlying distribution, we design an experiment under the same settings as before, but replace the background prior with a sample generated from an unknown noise process
into the model instead of $\bm{y}_{0}$. Here, $\mathcal{M}^{\mathrm{QG}}$ is the forward propagation model of the QG equations, and $\bm{x}_{-1}$ stands for the state at the previous time step. Comparisons between the step-wise RMSEs are displayed in fig. 13, where we fix $N=1$ and $p=0.25$. Despite the discrepancy in the background prior, our SOAD model demonstrates a capacity for self-correction within a few assimilation steps, exhibiting its robustness and relative insensitivity to inaccuracies in the background prior distribution.
Another important source of uncertainty in data assimilation arises from the nature of observational noise, which is often assumed to be Gaussian for analytical convenience. However, in reality, observational errors can exhibit non-Gaussian characteristics due to factors such as unknown nonlinear sensor transformations, mixed error sources, or limitations in the statistical characterization of measurement systems. To evaluate the robustness of our assimilation algorithm under such conditions, we perform experiments using a non-Gaussian noise model for the observations. In particular, compared to eq. 51, we set
$\bm{y}_{k}=\bm{S}_{k}\circ\mathcal{H}(\bm{x}_{k})+{\bm{\epsilon}_{k}},\quad{\bm{\epsilon}_{k}}\sim\mathscr{D},$ (58)
where the Gaussian noise assumption is replaced by a distribution $\mathscr{D}$, drawn from three types of element-wise noise models: Laplace distribution, Uniform distribution, and LogNormal distribution, whose density functions are visualized in fig. 14. Note that we have fixed appropriate parameters for each distribution so that all the distributions share the same mean and variance with our Gaussian assumption (51). The detailed parameters are listed in appendix C for reference.
Figure 14: The probabilistic density functions for various observational noise.
The quantitative results for assimilation under different observational noise distributions are summarized in tables 3 and 4. Each table reports the step-wise RMSE (scaled by $10^{-2}$) under various observation operators ($\mathcal{H}^{\mathrm{e}}$, $\mathcal{H}^{\mathrm{h}}$, and $\mathcal{H}^{\mathsf{v2v}}$), with and without the use of a background prior. The first table corresponds to the setting $N=1$, $p=1$, representing full spatio-temporal coverage, while the second table uses $N=2$, $p=0.25$, introducing a more difficult case. Across both settings, we observe that the performance of the SOAD model remains stable regardless of the specific noise distribution. In particular, the RMSE values under Laplace, Uniform, and moderately skewed LogNormal noise (e.g., LogNormal-0.1 to LogNormal-0.5) closely match those obtained under Gaussian noise, indicating that the model is robust to deviations from the Gaussian assumption. Only when the LogNormal noise becomes significantly skewed (LogNormal-1.0) do we see a slight degradation in performance, especially in the absence of a background prior. These results collectively suggest that SOAD is largely insensitive to the choice of noise distribution and maintains high assimilation accuracy even under non-Gaussian observational errors.
assimilation w/ prior assimilation w/o prior
[…]
Laplace $8.455^{\pm 0.46}$ $6.441^{\pm 0.36}$ $10.31^{\pm 0.49}$ $13.82^{\pm 0.53}$ $14.44^{\pm 1.2}$ $11.85^{\pm 0.76}$
Uniform $8.449^{\pm 0.46}$ $6.441^{\pm 0.36}$ $10.30^{\pm 0.49}$ $13.84^{\pm 0.53}$ $15.91^{\pm 5.6}$ $11.85^{\pm 0.76}$
LogNormal-0.1 $8.456^{\pm 0.47}$ $6.444^{\pm 0.36}$ $10.30^{\pm 0.49}$ $13.82^{\pm 0.53}$ $10.14^{\pm 3.5}$ $11.85^{\pm 0.75}$
LogNormal-0.2 $8.457^{\pm 0.47}$ $6.444^{\pm 0.36}$ $10.31^{\pm 0.49}$ $13.82^{\pm 0.53}$ $13.02^{\pm 5.0}$ $11.85^{\pm 0.75}$
LogNormal-0.5 $8.460^{\pm 0.47}$ $6.445^{\pm 0.36}$ $10.31^{\pm 0.49}$ $13.83^{\pm 0.53}$ $12.07^{\pm 6.1}$ $11.85^{\pm 0.75}$
LogNormal-1.0 $8.529^{\pm 0.48}$ $6.527^{\pm 0.37}$ $10.36^{\pm 0.49}$ $14.20^{\pm 0.65}$ $14.07^{\pm 1.6}$ $12.01^{\pm 0.78}$
Gaussian $\mathbf{8.457^{\pm 0.47}}$ $\mathbf{6.444^{\pm 0.36}}$ $\mathbf{10.30^{\pm 0.49}}$ $\mathbf{13.82^{\pm 0.53}}$ $\mathbf{12.93^{\pm 3.2}}$ $\mathbf{11.85^{\pm 0.75}}$
Table 3: RMSE (×$\times 10^{-2}$) with various observational noise distributions, $N=1$, $p=1$.
assimilation w/ prior assimilation w/o prior
[…]
Laplace $11.96^{\pm 1.1}$ $12.33^{\pm 1.3}$ $13.05^{\pm 1.2}$ $27.50^{\pm 1.0}$ $72.34^{\pm 9.3}$ $18.99^{\pm 0.80}$
Uniform $11.96^{\pm 1.1}$ $12.33^{\pm 1.3}$ $13.05^{\pm 1.2}$ $27.51^{\pm 1.0}$ $72.41^{\pm 9.4}$ $18.99^{\pm 0.80}$
LogNormal-0.1 $11.96^{\pm 1.1}$ $12.34^{\pm 1.3}$ $13.05^{\pm 1.2}$ $27.50^{\pm 1.0}$ $71.99^{\pm 9.2}$ $18.98^{\pm 0.80}$
LogNormal-0.2 $11.96^{\pm 1.1}$ $12.34^{\pm 1.3}$ $13.05^{\pm 1.2}$ $27.50^{\pm 1.0}$ $71.45^{\pm 8.4}$ $18.98^{\pm 0.81}$
LogNormal-0.5 $11.96^{\pm 1.1}$ $12.34^{\pm 1.3}$ $13.05^{\pm 1.2}$ $27.51^{\pm 1.0}$ $72.38^{\pm 9.0}$ $18.98^{\pm 0.81}$
LogNormal-1.0 $12.03^{\pm 1.1}$ $12.44^{\pm 1.3}$ $13.09^{\pm 1.2}$ $27.67^{\pm 1.0}$ $72.93^{\pm 7.4}$ $19.12^{\pm 0.84}$
Gaussian $\mathbf{11.96^{\pm 1.2}}$ $\mathbf{12.34^{\pm 1.3}}$ $\mathbf{13.05^{\pm 1.2}}$ $\mathbf{27.49^{\pm 0.53}}$ $\mathbf{72.14^{\pm 9.2}}$ $\mathbf{18.98^{\pm 0.80}}$
Table 4: RMSE (×$\times 10^{-2}$) with various observational noise distributions, $N=2$, $p=0.25$.